# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ 5.2 Inventory management
We consider a two-stage adjustable RO problem on inventory management (Iancu and Trichakis,, 2014, Section 5.2.1). There exist a retail network consisting of a single warehouse and $n$ different retail points, indexed by $i=1,\dots,n$. For simplicity, only a single product is sold. There are a total of $C$ units of the product available, and each retail point starts with 00 inventory and is capable of stocking at most $c_{i}$ units. The transportation costs for distributing the items at the $i$th retail point is $t_{i}$ currency units per unit of inventory, and the operating cost is $h_{i}$ per unit. The revenues per unit is $r_{i}$, which we take as our family parameter, ${\it i.e.},~{}y=r$, as these values are known but may be subject to minor changes across problem instances. Customer demand is uncertain, with a nominal demand plus an additional amount dependent on $m$ market factors. The formula for the demand at retail point $i$ is
$a_{i}=\bar{a}_{i}+q_{i}^{T}u,$
where $\bar{a}_{i}\in{\mbox{\bf R}}$ denotes the nominal demand, and $q_{i}\in{\mbox{\bf R}}^{m}$ denotes the exposure of demand $a_{i}$ to the market factors. We denote with $u$ the unknown market factors, which belong to an uncertainty set ${\mathcal{U}}(\theta)$. The decision variable is the stocking decisions for all retail points, denoted $s\in{\mbox{\bf R}}^{n}$. In addition, the problem involves a vector $w(u)$ of realized sales at each point, which is dependent on stocking decision $s$ and the unknown factors $u$ affecting the demand. We have
$w_{i}(u)=\min\{s_{i},a_{i}\},$
for each retail point. The manager makes stocking decisions to maximize worst-case profits, which is equivalent to minimizing worst-case loss, $-r^{T}w(u)+(t+h)^{T}s$, $\forall u\in{\mathcal{U}}(\theta)$. For each problem instance, the RO formulation is given as
@@
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}w(u)+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta)\\ &w_{i}(u)\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &w_{i}(u)\leq\bar{a}_{i}+q_{i}^{T}u,\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (17)
@@
where variables $s\in{\mbox{\bf R}}^{n}$, $w\in{\mbox{\bf R}}^{n}$, and $\tau\in{\mbox{\bf R}}$. To get a tractable approximation, we replace the adjustable decision $w_{i}(u)$ with their affine adjustable robust counterparts $z^{0}_{i}+v_{i}^{T}u$, where $z^{0}_{i}\in{\mbox{\bf R}}$ and $v_{i}\in{\mbox{\bf R}}^{f}$ are auxiliary introduced as part of the linear decision rules (Ben-Tal et al.,, 2009). The new formulation is
@@
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}z^{0}-r^{T}Vu+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta)\\ &z^{0}_{i}+v_{i}^{T}u\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &z^{0}_{i}+v_{i}^{T}u\leq\bar{a}_{i}+q_{i}^{T}u,\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (18)
@@
where $V=[v_{1}\cdots v_{n}]^{T}\in{\mbox{\bf R}}^{n\times f}$. We treat the multiple uncertain constraints as a single joint chance constraint represented as the maximum of concave functions.
@@
As in the work by Iancu and Trichakis, (2014), we let $n=10$. We consider the cases where the number of uncertain market factors is $m=4$ and $m=8$. We set $C=200$, and all other problem data are independently sampled from uniformly distributed random variables. Inventory capacities $c$ are between 30 and 50. Transportation and operation costs are between 1 and 3. Sales revenues $r$ are between 2 and 4, and nominal demand values are between 10 and 20. We consider 10 instances for $r$, perturbing each by random sample of a normal distribution with mean 0 and standard deviation $0.05$. The market factor exposure coefficients $q_{i}$ are between $-2$ and $2$. We generate uncertain market factors $u$ from 3 normal distributions, to simulate 3 uncertain market environments. For distribution $j$, the mean is $\mu_{i}=i\gamma_{j}$ for all $i=1,\dots,m$, with $\gamma=(0.3,0.6,0.9)$, and the variance is $\sigma_{i}=0.02^{2}+(0.1i)^{2}$ for all distributions. The total training sample size is $N=100$, split evenly across the 3 distributions. The testing sample size is $20000$.
Uncertainty set.
We train the shape and size of an ellipsoidal uncertainty set, with norm $p=2$, and for $\eta\in\eta^{\text{list}}=(0.01,0.03,0.05,0.08,0.10,0.15,0.2,0.3)$.
@@ Results.
In Table 4, we compare the performance of our LRO method with mean-variance RO and Wasserstein DRO, and note the desirable trade-off given by our learned reshaped set.
Table 4: Average inventory management performance over 20 runs. LRO: learning the shape and size of the uncertainty set with $\eta=0.3$. LRO-T0.03: Best value given by LRO with $\eta\in\eta^{\text{list}}$, tuned for an in-sample empirical probability of constraint violation ($\hat{\eta}^{\text{in}}$) below 0.03. LRO-T0.05: Best value given by LRO with $\eta\in\eta^{\text{list}}$, tuned for an $\hat{\eta}^{\text{in}}$ below 0.05. Mean-Var RO0.03: mean-variance RO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.03$. Mean-Var RO0.05: mean-variance RO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.05$. W-DRO0.03: Wasserstein DRO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.03$. W-DRO0.05: Wasserstein DRO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.05$. Obj.: the average out-of-sample objective value ${\mbox{\bf E}}[-r^{T}w(u)+(t+h)^{T}s]$. ${\rm time}$: average solve time after training.
[…]
In Figure 5, we show the trade-off between the average objective value and empirical probability of constraint violation for the 20 runs of the experiment. We observe an improved objective value for the LRO set across relevant levels of constraint violation. In Figure 5, for $m=4$, we note that the LRO set with a constraint violation $\hat{\eta}$ of around 0.028 achieves an objective value that the mean-variance RO and Wasserstein DRO sets only achieve at signficantly higher $\hat{\eta}$’s, of around 0.12 and 0.07, respectively. A similar comparison holds for $m=8$. It is apparent that our decision-focused manner of learning uncertainty sets identifies more effective regions of uncertainty and therefore reduces the conservatism of the data-driven solutions.
Figure 5: Inventory. Average trade-offs between out-of-sample objective values and probability of constraint violation across 20 experiments. The shaded region represents the 0.3 to 0.7 quantiles. The LROPT results are obtained through training procedures with $\eta$ values in the range $[0,0.3]$.
@@ 6 Conclusion
We develop a data-driven method to learn the shape and size uncertainty sets in robust optimization in a decision-focused manner, i.e., using the structure and the solution of the robust counterpart. By formulating the learning problem as a stochastic optimization problem with $\mathop{\bf CVaR}$ constraints, we choose the shape and size of the uncertainty sets to optimize the data-driven performance while enforcing desired probabilistic guarantees. We show convergence of our algorithm using nonsmooth implicit differentiation, and provide finite sample probabilistic guarantees on unseen data. Through numerical examples, we demonstrate the performance gain achieved by our method compared to traditional manual tuning methods.
@@ 7 Acknowledgments
Irina Wang and Bartolomeo Stellato are supported by the NSF CAREER Award ECCS 2239771. The simulations presented in this article were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University.
@@ Appendix A Uncertainty sets and general reformulations
We derive convex reformulations of (1) for a general class of constraints $g$ and common uncertainty sets. We express $u$ as an affine transformation of an uncertain parameter $z$, i.e., $u=Ax+b$, where $A\in{\mbox{\bf R}}^{m\times\hat{m}}$, and $\hat{m}\leq m$, and $b\in{\mbox{\bf R}}^{m}$. The proofs are in Appendix E. For notational simplicity, we suppress the dependency of $g$ on $y$ here and for all following sections on reformulations. We consider two general types of functions $g$: the maximum of concave functions (left), where each $g_{l}$ is concave in $u$, and its special case, the maximum of affine functions (right), where $P_{l}(x)$ and $a_{l}(x)$ are convex functions in $x$:
$g(x,u)=\max_{l=1,\dots,L}g_{l}(x,u),\qquad g(x,u)=\max_{l=1,\dots,L}(P_{l}(x)u+a_{l}(x)).$
@@
${\mathcal{U}}(\theta)=\left\{u=Az+b\mid\|z\|_{p}\leq 1,\quad u\in S\right\},$
where $A\in{\mbox{\bf R}}^{m\times\hat{m}}$, $b\in{\mbox{\bf R}}^{m}$, $p\in{\mbox{\bf R}}$ , and $S$ includes additional support information such as upper and lower bounds. We let $p$ be the order of the norm, which is any integer value $\geq 1$, or equal to $\infty$. In the latter case, this becomes the box uncertainty set. The trainable parameter $\theta=(A,b)$. Note that if $A=\rho I$, $b=0$, and $S={\mbox{\bf R}}^{m}$, this is equivalent to the traditional ellipsoidal uncertainty set ${\{u\mid\|u\|_{p}\leq\rho\}}$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}\text{minimize}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|A^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L,\end{array}$
with auxiliary variables $\omega_{l},h_{l}\in{\mbox{\bf R}}^{m}$ for $l=1,\dots,L$, and $g^{\star}(x,h)$ the conjugate function. This result follows from the theorem of infimal convolutions and composites for conjugate functions (Rockafellar and Wets,, 1998)(Bauschke and Combettes,, 2017). As a special case, we have the reformulation for the maximum of affine constraint,
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega_{l})+\|A^{T}(\omega_{l}-P_{l}(x))\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
@@
The budget uncertainty set requires both box and ellipsoidal constraints, and is defined as
${\mathcal{U}}_{\text{bud}}(\theta)=\{u=Az+b\mid\|z\|_{\infty}\leq\rho_{1},\quad\|z\|_{1}\leq\rho_{2},\quad u\in S\},$
where $\theta=(A,b,\rho_{1},\rho_{2})$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L,\end{array}$
with $r_{l}\in{\mbox{\bf R}}^{\hat{m}}$, $\omega_{l},h_{l}\in{\mbox{\bf R}}^{m}$ for $l=1,\dots,L$ as auxiliary variables. The maximum of affine reformulation is
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega_{l})+\rho_{1}\|r_{l}-A^{T}(P_{l}(x)+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L.\end{array}$ (19)
@@
${\mathcal{U}}(\theta)=\{u=Az+b\mid Dz\leq d,\quad u\in S\},$
where $\theta=(D,d)\in{\mbox{\bf R}}^{r\times\hat{m}}\times{\mbox{\bf R}}^{r}$ defines the polyhedron. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
with auxiliary variables $\gamma_{l},\omega_{l}\in{\mbox{\bf R}}^{m}$, $h_{l}\in{\mbox{\bf R}}^{r},$ for $l=1,\dots,L$. For the maximum of affine constraint, we have:
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+d^{T}h_{l}+b^{T}(P_{l}(x)-\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(P_{l}(x)-\gamma_{l})=D^{T}h_{l},\quad l=1,\dots,L\\ &h_{l}\geq 0,\quad l=1,\dots,L.\end{array}$

### v5

@@ 6.2 Inventory management
We consider a two-stage adjustable RO problem on inventory management (Iancu and Trichakis,, 2014, Section 5.2.1). There exist a retail network consisting of a single warehouse and $n$ different retail points, indexed by $i=1,\dots,n$. For simplicity, only a single product is sold. There are a total of $C$ units of the product available, and each retail point starts with 00 inventory and is capable of stocking at most $c_{i}$ units. The transportation costs for distributing the items at the $i$th retail point is $t_{i}$ currency units per unit of inventory, and the operating cost is $h_{i}$ per unit. The revenues per unit is $r_{i}$, which we take as a context parameter, as these values are known but may be subject to minor changes across problem instances. Customer demand $u\in{\mathbf{R}}^{n}$ is uncertain, but we have access to contextual data on $q$ market factors, $F\in{\mathbf{R}}^{(n,q)}$, which influences the demand covariance. The total contextual information is thus $x=(r,F)$. The decision variable is the stocking decisions for all retail points, denoted $s\in{\mathbf{R}}^{n}$. In addition, the problem involves a vector $w(u)$ of realized sales at each point, which is dependent on stocking decision $s$ and the unknown factors $u$ affecting the demand; we have $w_{i}(u)=\min\{s_{i},u_{i}\}$ for each retail point. The manager makes stocking decisions to maximize worst-case profits, which is equivalent to minimizing worst-case loss, $-r^{T}w(u)+(t+h)^{T}s$, $\forall u\in{\mathcal{U}}(\theta,x)$. For each problem instance, the RO formulation is given as
@@
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}w(u)+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &w_{i}(u)\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &w_{i}(u)\leq u_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (19)
@@
where variables $s\in{\mathbf{R}}^{n}$, $w\in{\mathbf{R}}^{n}$, and $\tau\in{\mathbf{R}}$. To get a tractable approximation, we replace the adjustable decision $w_{i}(u)$ with their affine adjustable robust counterparts $z^{0}_{i}+v_{i}^{T}u$, where $z^{0}_{i}\in{\mathbf{R}}$ and $v_{i}\in{\mathbf{R}}^{f}$ are auxiliary introduced as part of the linear decision rules (Ben-Tal et al.,, 2009). The new formulation is
@@
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}z^{0}-r^{T}Vu+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &z^{0}_{i}+v_{i}^{T}u\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &z^{0}_{i}+v_{i}^{T}u\leq u_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (20)
@@
where $V=[v_{1}\cdots v_{n}]^{T}\in{\mathbf{R}}^{n\times f}$. We treat the $12n+$ 1 uncertain constraints as a single joint chance constraint represented as the maximum of concave functions.
@@
As in the work by Iancu and Trichakis, (2014), we let $n=10$. This is also the size of the uncertain parameter, i.e., $m=n$. We consider the case where the number of market factors for the contextual information is $q=4$. We set $C=200$, and all other problem data are independently sampled from uniformly distributed random variables. Inventory capacities $c$ are between 30 and 50. Transportation and operation costs are between $0.1$ and $0.3$, and sales revenues $r$ are between 2 and 4. We consider 10 instances for $r$, perturbing each by random sample of a normal distribution with mean 0 and standard deviation $0.05$. We also consider 10 instances for $F$, each drawn from a standard $N(0,1)^{(n,q)}$ distribution. The contextual uncertain demand $u^{i}$ is sampled from a multivariate normal distribution $N(\mu^{i}(r^{i})),\Sigma^{i}(F^{i}))$, where $\mu^{i}(r^{i})=\mu-0.1r^{i}$, and $\Sigma^{i}(F^{i})=0.3F^{i}F^{iT}$. The nominal demand $\mu$ is sampled uniformly on $[10,20]$. We consider the case with $N=1000$ samples, with 100 samples for each context. We use $30\%$ of the data for training, $20\%$ for validation, and $50\%$ for testing.
Uncertainty sets.
As this problem has $212n+1=$ 21 uncertain constraints, the ECRO uncertainty set, which works with uncertain objectives only, does not apply. In addition, the LCX uncertainty set would require a separate cutting-plane procedure for each uncertain constraint, and also for each context, so we opted to forgo it. For the non-contextual Wasserstein DRO method (W), we use the mean robust optimization (MRO) approximation, with the number of clusters $K=30$, given by Wang et al., (2024), to obtain a more tractable formulation. For our training method, due to the higher complexity of the problem, we lower the batch size to 30, and the initial step size to $0.0005$. We halve the step size every $20$ steps. To adjust for the smaller step sizes, we increase the number of outer iterations to $25$, bringing the total number of iterations to $500$.
@@ Results.
In Table 5, we compare the performance of all methods, and note that our method again outperforms the others in all objective metrics. Our decision-focused manner of learning uncertainty sets identifies more effective regions of uncertainty, and therefore reduces the conservatism of the data-driven solutions.
Table 5: Average inventory management performance over 10 runs. The best value for each category is in bold.
$n=10$ $N=1000$ Method $\hat{\eta}$ Worst-case Average-case $\mathop{\bf CVaR}$ LRO $\bf{0.0958}$ $\bf{-401.59}$ $\bf{-404.48}$ $\bf{-401.49}$ MV $0.1064$ $-385.71$ $-394.01$ $-385.71$ CMV $0.0986$ $-378.07$ $-397.78$ $-378.07$ W $0.0994$ $-376.06$ $-383.44$ $-375.47$ CW $0.0998$ $-388.94$ $-396.21$ $-387.97$ SA $0.4597$ $-395.37$ $-395.51$ $-395.32$
@@ 7 Conclusion
We develop a data-driven method to learn the contextual shape of uncertainty sets in a decision-focused manner, i.e., using the structure and the solution of the robust counterpart. By formulating the learning problem as a stochastic optimization problem with $\mathop{\bf CVaR}$ constraints, we can optimize over a weighted performance metric while enforcing desired probabilistic guarantees. We show convergence of our algorithm using nonsmooth implicit differentiation, and provide finite sample probabilistic guarantees on unseen data. Through numerical examples, we demonstrate the performance gain achieved by our method compared to traditional manual tuning methods, for both worst-case and average-case performance.
@@ 8 Acknowledgments
Irina Wang and Bartolomeo Stellato are supported by the NSF CAREER Award ECCS 2239771. The simulations presented in this article were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University. We would like to thank Jason Klusowski for the useful feedback and discussions on empirical process theory, and Cole Becker for his contributions to the original ideation of this work.
@@ Appendix A Uncertainty sets and general reformulations
We derive convex reformulations of (1) for a general class of constraints $g$ and common uncertainty sets. We express $u$ as an affine transformation of an uncertain parameter $v$, i.e., $u=A(x)v+b(x)$, where $A(x)\in{\mathbf{R}}^{m\times\hat{m}}$, and $\hat{m}\leq m$, and $b(x)\in{\mathbf{R}}^{m}$. The proofs are in Appendix E. For notational simplicity, we suppress the dependency of $g$ on $x$ here and for all following sections on reformulations. We consider two types of functions $g$: the maximum of concave functions (left), where each $g_{l}$ is concave in $u$, and its special case, the maximum of affine functions (right), where $P_{l}(z)$ and $a_{l}(z)$ are convex functions in $z$:
$g(z,u)=\max_{l=1,\dots,L}g_{l}(z,u),\qquad g(z,u)=\max_{l=1,\dots,L}(P_{l}(z)u+a_{l}(z)).$
@@
${\mathcal{U}}(\theta,x)=\left\{u=A(x)v+b(x)\mid\|v\|_{p}\leq 1,\quad u\in S\right\},$
where $A(x)\in{\mathbf{R}}^{m\times\hat{m}}$, $b(x)\in{\mathbf{R}}^{m}$, $p\in{\mathbf{R}}$ , and $S$ includes additional support information such as upper and lower bounds. We let $p$ be the order of the norm, which is any integer value $\geq 1$, or equal to $\infty$. In the latter case, this becomes the box uncertainty set. The trainable parameter $\theta=(A(\cdot),b(\cdot))$, the mapping from the context parameter to the reshaping matrices. Note that if $A=\rho I$, $b=0$, and $S={\mathbf{R}}^{m}$, this is equivalent to the traditional ellipsoidal uncertainty set ${\{u\mid\|u\|_{p}\leq\rho\}}$. For the maximum of concave constraint, we have the following reformulation,
$\begin{array}[]{ll}\text{minimize}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{l}+\omega_{l})+\|A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L,\end{array}$
with auxiliary variables $\omega_{l},h_{l}\in{\mathbf{R}}^{m}$ for $l=1,\dots,L$, and $g^{\star}(z,h)$ the conjugate function. This result follows from the theorem of infimal convolutions and composites for conjugate functions (Rockafellar and Wets,, 1998)(Bauschke and Combettes,, 2017). In the special (maximum of affine) case, we instead have
$\begin{array}[]{ll}\mbox{minimize}&f(z)\\ \mbox{subject to}&a_{l}(z)+\sigma_{S}(\omega_{l})+b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(P_{l}(z)-\omega_{l})+\|A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{l}-P_{l}(z))\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
@@
The budget uncertainty set requires both box and ellipsoidal constraints,
${\mathcal{U}}_{\text{bud}}(\theta,x)=\{u=A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}v+b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}\mid\|v\|_{\infty}\leq\rho_{1},\quad\|v\|_{1}\leq\rho_{2},\quad u\in S\},$
where $\theta=(A(\cdot),b(\cdot),\rho_{1},\rho_{2})$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L,\end{array}$
with $r_{l}\in{\mathbf{R}}^{\hat{m}}$, $\omega_{l},h_{l}\in{\mathbf{R}}^{m}$ for $l=1,\dots,L$ as auxiliary variables. The maximum of affine reformulation is
$\begin{array}[]{ll}\mbox{minimize}&f(z)\\ \mbox{subject to}&a_{l}(z)+\sigma_{S}(\omega_{l})+b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(P_{l}(z)-\omega_{l})+\rho_{1}\|r_{l}-A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(P_{l}(z)+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L.\end{array}$ (21)
@@
${\mathcal{U}}(\theta)=\{u=A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}v+b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}\mid Dv\leq d,\quad u\in S\},$
where $\theta=(A(\cdot),b(\cdot),D,d)$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,\omega_{l})-b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
with auxiliary variables $\gamma_{l},\omega_{l}\in{\mathbf{R}}^{m}$, $h_{l}\in{\mathbf{R}}^{r},$ for $l=1,\dots,L$. The maximum of affine reformulation is
$\begin{array}[]{ll}\mbox{minimize}&f(z)\\ \mbox{subject to}&a_{l}(z)+d^{T}h_{l}+b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(P_{l}(z)-\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(P_{l}(z)-\gamma_{l})=D^{T}h_{l},\quad l=1,\dots,L\\ &h_{l}\geq 0,\quad l=1,\dots,L.\end{array}$